OpenStudy (anonymous):

Find the product: (x + 2)( x − 2)

OpenStudy (anonymous):

You simply multiply

OpenStudy (anonymous):

Here: (x+2) times (x-2). You don't know?

OpenStudy (anonymous):

Use the F.O.I.L Method. Multiply: First, outside, inside, last First: x*x= x^2 Outside: x*(-2)= -2x Inside: 2*x= 2x Last; 2*(-2)= -4 Then add all the terms: x^2 + (-2x) + 2x + (-4) = x^2-4

OpenStudy (anonymous):

formula\[(a-b)(a+b) = a ^{2} - b ^{2}\] \[(x-2)(x+2) = x ^{2}\]

OpenStudy (anonymous):

sorry \[x ^{2} - 4 \]